# Nonlinearly Preconditioned Gradient Methods under Generalized Smoothness

### v1

@@ Appendix D Details on the second-order condition
In this subsection we examine the second-order characterizations described in Lemma 2.5 for the isotropic reference functions $\phi_{i}$ generated by $h_{1}(x)=\cosh(x)-1$, $h_{2}(x)=\exp(|x|)-|x|-1$, $h_{3}(x)=-|x|-\ln(1-|x|)$. As already mentioned in the introduction, in this case we have
@@
and $\nabla^{2}\phi_{i}^{*}(y)={h_{i}^{*}}^{\prime\prime}(\|y\|)I$ otherwise. For ease of presentation we denote $\alpha_{i}(y)={h_{i}^{*}}^{\prime\prime}(\|y\|)$ and $\beta_{i}(y)=\frac{{h_{i}^{*}}^{\prime}(\|y\|)}{\|y\|}$.
D.1 Norm to power
We begin our examples with $f(x)=\tfrac{1}{p}\|x\|^{p}$ for $p\geq 4$. The gradient and Hessian of $f$ are given respectively by
@@ D.1 Norm to power
$\nabla^{2}\phi_{i}^{*}(\bar{L}^{-1}\nabla f(x))$ $=\alpha_{i}(\bar{L}^{-1}\|\nabla f(x)\|)\frac{\nabla f(x)\nabla f(x)^{\top}}{\|\nabla f(x)\|^{2}}+\beta_{i}(\bar{L}^{-1}\|\nabla f(x)\|)\left(I-\frac{\nabla f(x)\nabla f(x)^{\top}}{\|\nabla f(x)\|^{2}}\right)$
$=\alpha_{i}(\bar{L}^{-1}\|x\|^{p-1})\frac{xx^{\top}}{\|x\|^{2}}+\beta_{i}(\bar{L}^{-1}\|x\|^{p-1})\left(I-\frac{xx^{\top}}{\|x\|^{2}}\right).$
@@
$\nabla^{2}\phi_{i}^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x)=(p-1)\alpha_{i}(\bar{L}^{-1}\|x\|^{p-1})\|x\|^{p-4}xx^{\top}+\beta_{i}(\bar{L}^{-1}\|x\|^{p-1})\|x\|^{p-2}\left(I-\frac{xx^{\top}}{\|x\|^{2}}\right).$
The largest eigenvalue of this matrix is
$\lambda_{\max}(\nabla^{2}\phi_{i}^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x))=\max\Big{\{}(p-1)\alpha_{i}(\bar{L}^{-1}\|x\|^{p-1}),\beta_{i}(\bar{L}^{-1}\|x\|^{p-1})\Big{\}}\|x\|^{p-2}$ (30)
and in all of the cases considered in this subsection, $\max(p-1)\alpha_{i}(\bar{L}^{-1}\|x\|^{p-1})\geq\max\beta_{i}(\bar{L}^{-1}\|x\|^{p-1})$. Therefore, the inequality of Definition 2.4 dictates $(p-1)\alpha_{i}(\bar{L}^{-1}\|x\|^{p-1})\|x\|^{p-2}\leq L\bar{L}$ for all $x\in\mathbb{R}^{n}$. We take $p=4$ in order to make the presentation clearer. We now specialize these results to the reference functions we consider in this section:
  • For $h_{1}(x)=\cosh(x)-1$, we obtain $\frac{3\|x\|^{2}}{\sqrt{1+\bar{L}^{-2}\|x\|^{6}}}\leq L\bar{L}$ or equivalently $\frac{3\|x\|^{2}}{\sqrt{\bar{L}^{2}+\|x\|^{6}}}\leq L$. The quantity on the l.h.s. is maximized at $\|x\|=(2\bar{L}^{2})^{1/6}$ and thus we obtain that for any $\bar{L}>0$, $L\geq(2\bar{L}^{-1})^{1/3}\sqrt{3}$.
  • For $h_{2}(x)=\exp(|x|)-|x|-1$, we have $\frac{3\|x\|^{2}}{\bar{L}+\|x\|^{3}}\leq L$. The quantity on the l.h.s. is maximized at $\|x\|=(2\bar{L})^{1/3}$ and thus we have for $\bar{L}>0$, $L\geq\frac{2^{2/3}}{\bar{L}^{1/3}}$.
  • For $h_{3}(x)=-|x|-\ln(1-|x|)$, we obtain $\frac{3\bar{L}\|x\|^{2}}{(\bar{L}+\|x\|^{3})^{2}}\leq L$. Now, the quantity on the l.h.s. is maximized at $\|x\|=(\frac{\bar{L}}{2})^{1/3}$ and thus, for $\bar{L}>0$, $L\geq\frac{2^{4/3}}{3\bar{L}^{1/3}}$.
D.2 Logistic loss
We now consider the logistic loss function $f(x)=\log(1+\exp(-\alpha^{\top}x))$. The gradient and the hessian of $f$ are given respectively by
@@ D.2 Logistic loss
${h^{*}}^{\prime\prime}(\bar{L}^{-1}\|\nabla f(x)\|)\frac{\|\alpha\|^{2}}{\exp(-\alpha^{\top}x)(1+\exp(\alpha^{\top}x))^{2}}\leq L\bar{L}.$ (32)
  • For $h_{1}(x)=\cosh(x)-1$, we have $\frac{\|\alpha\|^{2}}{(1+\exp(-\alpha^{\top}x))\sqrt{\bar{L}^{2}(1+\exp(\alpha^{\top}x))+\|\alpha\|^{2}}}\leq\frac{2\|\alpha\|^{2}}{\sqrt{27\bar{L}^{2}+4\|\alpha\|^{2}}}\leq L$.
  • For $h_{2}(x)=\exp(|x|)-|x|-1$, we obtain $\frac{\|\alpha\|^{2}}{(\bar{L}(1+\exp(\alpha^{\top}x))+\|a\|)(1+\exp(\alpha^{\top}x))}\leq\frac{\|\alpha\|^{2}}{4\bar{L}+\|\alpha\|}\leq L$.
  • For $h_{3}(x)=-|x|-\ln(1-|x|)$, we have $\frac{\bar{L}\|\alpha\|^{2}}{(\bar{L}(1+\exp(\alpha^{\top}x))+\|\alpha\|)^{2}\exp(-\alpha^{\top}x)}\leq\frac{\bar{L}\|\alpha\|^{2}+\|\alpha\|^{3}}{4(\bar{L}+\|\alpha\|)^{2}}\leq L.$
D.3 Clipping
@@
We now consider the case that $\phi=\frac{1}{2}\|\cdot\|^{2}+\delta_{B}(\cdot)$ where $B$ is the unit norm ball. By [Themelis2019, p. 404], the generalized Jacobian of the preconditioner is given by
$\partial_{C}(\nabla\phi^{*})(x)=\begin{cases}\{I\},&\textnormal{if $\|x\|<1$,}\\ \operatorname{con}\{I,\|x\|^{-1}(I-\overline{\operatorname{sgn}}(x)\overline{\operatorname{sgn}}(x)^{\top}\},&\textnormal{if $\|x\|=1$,}\\ \{\|x\|^{-1}(I-\overline{\operatorname{sgn}}(x)\overline{\operatorname{sgn}}(x)^{\top})\},&\textnormal{if $\|x\|>1$.}\end{cases}$
The second-order condition then becomes
@@
$\begin{cases}\lambda_{\rm max}(\nabla^{2}f(x))\leq L\bar{L}I,&\textnormal{if $\|\nabla f(x)\|<\bar{L}$,}\\ \lambda_{\rm max}\left(\lambda\nabla^{2}f(x)+(1-\lambda)\bar{L}\|\nabla f(x)\|^{-1}\Pi_{\nabla f(x)^{\perp}}\nabla^{2}f(x)\right)\leq L\bar{L},&\textnormal{if $\|\nabla f(x)\|=\bar{L}$,}\\ \lambda_{\rm max}(\Pi_{\nabla f(x)^{\perp}}\nabla^{2}f(x))\leq L\|\nabla f(x)\|,&\textnormal{if $\|\nabla f(x)\|>\bar{L}$,}\end{cases}$
for all $\lambda\in[0,1]$, with $\Pi_{\nabla f(x)^{\perp}}$ denoting the projection matrix onto the orthogonal complement of the subspace spanned by $\nabla f(x)$. If $\nabla^{2}f(x)=C(x)\nabla f(x)\nabla f(x)^{\top}$ for some $C:\mathbb{R}^{n}\to\mathbb{R}$, then the last condition is vacuous, and we only require that $\nabla^{2}f(x)\preceq L\bar{L}I$ for $\|\nabla f(x)\|\leq\bar{L}$. Otherwise, we require that
$\begin{cases}\nabla^{2}f(x)\preceq L\bar{L}I,&\textnormal{if $\|\nabla f(x)\|\leq\bar{L}$,}\\ \nabla^{2}f(x)\preceq L\|\nabla f(x)\|,&\textnormal{if $\|\nabla f(x)\|\geq\bar{L}$.}\end{cases}$
Note that these conditions can be interpreted as $(L_{0},L_{1})$-smoothness with $L_{1}=0$ and $L_{0}=0$ respectively.

### v2

@@ Appendix D Details on the second-order condition
In this section we provide further details on the second-order condition Definition 2.4. We complement the discussion in Section 2 by showing that the norm-coercivity condition on the forward operator in Definition 2.4 is in fact mild even for functions $\phi$ with full domain.
Proposition D.1.
Let $\phi=h\circ\|\cdot\|$ such that $h\in\mathcal{C}^{1}(\mathbb{R})$ satisfies Assumption 1.1. If there exists some $C>0$ such that
$\|\nabla f(x)\|\leq\bar{L}|h^{\prime}(L\|x\|)|$ (31)
for all $x$ such that $\|x\|\geq C$, then
$\lim_{\|x\|\to\infty}\|T_{{\delta L^{-1}},{\bar{L}^{-1}}}(x)\|=\infty,$
for all $\delta<1$.
Proof.
In the following we assume that $\|x\|$ is large enough such that the assumption of the proposition holds. We have that
$\bar{L}^{-1}\|\nabla f(x)\|\leq|h^{\prime}(L\|x\|)|\Rightarrow{h^{*}}^{\prime}(\bar{L}^{-1}\|\nabla f(x)\|)\leq L\|x\|.$ (32)
The implication follows since $h(0)\geq h(t)-h^{\prime}(t)t$, meaning that $h^{\prime}(t)\geq 0$ for $t\geq 0$ and thus $|h^{\prime}(t)|=h^{\prime}(t)$ on this interval implying ${h^{*}}^{\prime}(|h^{\prime}(t)|)=t$. Now, by the reverse triangle inequality:
$\|T_{{\delta L^{-1}},{\bar{L}^{-1}}}(x)\|$ $\geq\|x\|-\delta L^{-1}|{h^{*}}^{\prime}(\bar{L}^{-1}\|\nabla f(x)\|)|$
$\geq\|x\|(1-\delta),$
where the second inequality follows by (32). Therefore, since $\delta<1$, $\lim_{\|x\|\to\infty}\|T_{{\delta L^{-1}},{\bar{L}^{-1}}}(x)\|=\infty$ and the proof is complete. ∎
The fact that this condition is quite mild can be seen by choosing $\phi(x)=\cosh(\|x\|)-1$, where we allow $\|\nabla f(x)\|$ to grow exponentially with $\|x\|$. It is straightforward that this condition holds for example when the norm of the gradient is bounded by some polynomial of the norm of $x$ when $\|x\|$ is large enough.
We next show that when the matrix $H\nabla^{2}f(x)$ is symmetric, the norm-coercivity property of the forward operator in Definition 2.4 is not required in order to obtain a result similar to Proposition 2.6.
Proposition D.2.
Let $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ be such that for all $x\in\mathbb{R}^{n}$ and $H\in\partial_{C}(\nabla\phi^{*})(\bar{L}^{-1}\nabla f(x))$,
$\lambda_{\rm\max}(H\nabla^{2}f(x))\leq L\bar{L}$ (33)
and $H\nabla^{2}f(x)$ is symmetric. Then, the following inequality holds:
$\langle T_{{\gamma},{\bar{L}^{-1}}}(x)-T_{{\gamma},{\bar{L}^{-1}}}(\bar{x}),x-\bar{x}\rangle\geq(1-\gamma L)\|x-\bar{x}\|^{2},$ (34)
for all $x,\bar{x}\in\mathbb{R}^{n}$. In particular, for $\gamma<L^{-1}$, the forward operator is strongly monotone with parameter $1-\gamma L$ and thus injective.
Proof.
Note that the mapping $\nabla\phi^{*}\circ(\bar{L}^{-1}\nabla f)$ is locally Lipschitz, since $\nabla\phi^{*}$ is globally Lipschitz and $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$. Then, we can invoke the generalized mean value theorem in its summation form (Facchinei & Pang, 2003, Proposition 7.1.16): for two points $x,\bar{x}\in\mathbb{R}^{n}$, there exist $n$ points $z_{i}\in(x,\bar{x})$ and $n$ scalars $\alpha_{i}\geq 0$ summing to unity such that
where $V_{i}\in\partial_{C}(\nabla\phi^{*}\ \circ\bar{L}^{-1}\nabla f)(z_{i})$. Now, in light of (Clarke, 1990, p. 75),
$\partial_{C}(\nabla\phi^{*}\circ\bar{L}^{-1}\nabla f)(x)v\subseteq\operatorname{con}\{\partial_{C}(\nabla\phi^{*})(\bar{L}^{-1}\nabla f(x))\bar{L}^{-1}\nabla^{2}f(x)v\}$
for any $v\in\mathbb{R}^{n}$. Therefore, any $V_{i}(x-\bar{x})$ can be written as $\bar{L}^{-1}\sum_{j=1}^{d}\beta_{j}H_{j}\nabla^{2}f(z_{i})(x-\bar{x})$, for $d>0$, with $H_{j}\in\partial_{C}(\nabla\phi^{*})(\bar{L}^{-1}\nabla f(z_{i}))$ and $\beta_{j}\geq 0$ summing to unity.
Taking an inner product with $(x-\bar{x})$, we have that
$\langle\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x))-\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x})),x-\bar{x}\rangle$ $=\bar{L}^{-1}\sum_{i=1}^{n}\alpha_{i}\sum_{j=1}^{d}\beta_{j}\langle x-\bar{x},H_{j}\nabla^{2}f(z_{i})(x-\bar{x})\rangle$
$\leq\bar{L}^{-1}\sum_{i=1}^{n}\alpha_{i}\bar{L}L\|x-\bar{x}\|^{2}$
$=L\|x-\bar{x}\|^{2},$
implying that
which is the claimed result. ∎
D.1 Examples
We now move on to providing examples of functions satisfying Definition 2.4. We consider the reference functions $\phi_{1}(x)=\cosh(\|x\|)-1$, $\phi_{2}(x)=\exp(\|x\|)-\|x\|-1$ and $\phi_{3}(x)=-\|x\|-\ln(1-\|x\|)$, which are generated by the ($1$-dimensional) kernel functions $h_{1}(x)=\cosh(x)-1$, $h_{2}(x)=\exp(|x|)-|x|-1$ and $h_{3}(x)=-|x|-\ln(1-|x|)$.
We first recall the preconditioner $\nabla\phi^{*}$ and its Jacobian $\nabla^{2}\phi^{*}$ for general isotropic reference functions.
@@ D.1 Examples
and $\nabla^{2}\phi_{i}^{*}(y)={h_{i}^{*}}^{\prime\prime}(\|y\|)I$ otherwise. For ease of presentation we denote $a_{i}(y)={h_{i}^{*}}^{\prime\prime}(\|y\|)$ and $b_{i}(y)=\frac{{h_{i}^{*}}^{\prime}(\|y\|)}{\|y\|}$.
Table 2: Anisotropic smoothness constants for the examples of subsection D.1
. $\phi_{1}$ $\phi_{2}$ $\phi_{3}$ $L_{\rm norm}$ $\frac{2^{1/3}\sqrt{3}}{\bar{L}^{1/3}}$ $\frac{2^{2/3}}{\bar{L}^{1/3}}$ $\frac{2^{4/3}}{3\bar{L}^{1/3}}$ $L_{\rm logistic}$ $\frac{\|\alpha\|^{2}}{\sqrt{16\bar{L}^{2}+\|\alpha\|^{2}}}$ $\frac{\|\alpha\|^{2}}{4\bar{L}+\|\alpha\|}$ $\frac{\bar{L}\|\alpha\|^{2}+\|\alpha\|^{3}}{4(\bar{L}+\|\alpha\|)^{2}}$
Example D.3 (Norm to power).
Let $f(x)=\tfrac{1}{4}\|x\|^{4}$. Then, $f$ is $(L,\bar{L})$-anisotropically smooth relative to $\phi_{i}$ for any $\bar{L}>0$ and $L>L_{\rm norm}$ from the first row of Table 2.
Proof.
We first consider the more general $f(x)=\frac{1}{p}\|x\|^{p}$ with $p\geq 4$. The gradient and Hessian of $f$ are given respectively by
@@
$\nabla^{2}\phi_{i}^{*}(\bar{L}^{-1}\nabla f(x))$ $=a_{i}(\bar{L}^{-1}\|\nabla f(x)\|)\frac{\nabla f(x)\nabla f(x)^{\top}}{\|\nabla f(x)\|^{2}}+b_{i}(\bar{L}^{-1}\|\nabla f(x)\|)\left(I-\frac{\nabla f(x)\nabla f(x)^{\top}}{\|\nabla f(x)\|^{2}}\right)$
$=a_{i}(\bar{L}^{-1}\|x\|^{p-1})\frac{xx^{\top}}{\|x\|^{2}}+b_{i}(\bar{L}^{-1}\|x\|^{p-1})\left(I-\frac{xx^{\top}}{\|x\|^{2}}\right).$
@@
$\nabla^{2}\phi_{i}^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x)=(p-1)a_{i}(\bar{L}^{-1}\|x\|^{p-1})\|x\|^{p-2}\frac{xx^{\top}}{\|x\|^{2}}+b_{i}(\bar{L}^{-1}\|x\|^{p-1})\|x\|^{p-2}\left(I-\frac{xx^{\top}}{\|x\|^{2}}\right).$
The largest eigenvalue of this (symmetric) matrix is
$\lambda_{\max}(\nabla^{2}\phi_{i}^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x))=\max\Big{\{}(p-1)a_{i}(\bar{L}^{-1}\|x\|^{p-1}),b_{i}(\bar{L}^{-1}\|x\|^{p-1})\Big{\}}\|x\|^{p-2}.$ (37)
and for all the reference functions we consider in this subsection, $(p-1)a_{i}(\bar{L}^{-1}\|x\|^{p-1})\geq b_{i}(\bar{L}^{-1}\|x\|^{p-1})$. Therefore the inequality (5) dictates $(p-1)a_{i}(\bar{L}^{-1}\|x\|^{p-1})\|x\|^{p-2}<L\bar{L}$ for all $x\in\mathbb{R}^{n}$.
Now we specialize to each of the reference functions we consider as well as take $p=4$.
For $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$, we obtain the conditions
$\phi_{1}:\frac{3\|x\|^{2}}{\sqrt{\bar{L}^{2}+\|x\|^{6}}}<L,\qquad\phi_{2}:\frac{3\|x\|^{2}}{\bar{L}+\|x\|^{3}}<L,\qquad\phi_{3}:\frac{3\bar{L}\|x\|^{2}}{(\bar{L}+\|x\|^{3})^{2}}<L$
for which the left-hand sides are maximized at $\|x\|=(2\bar{L}^{2})^{1/6}$, $\|x\|=(2\bar{L})^{1/3}$ and $\|x\|=(\frac{\bar{L}}{2})^{1/3}$ respectively. Plugging these values in yields the resulting lower bounds $L_{\rm norm}$ in Table 2.
Since in every case $H\nabla^{2}f(x)$ is a symmetric matrix, it follows from Proposition D.2 that the operator $T_{{\delta L^{-1}},{\bar{L}^{-1}}}$ is injective for any $\delta<1$. This implies the anisotropic smoothness of $f$ relative to $\phi_{3}$ in light of Proposition 2.9 since $\operatorname{dom}\phi_{3}$ is bounded. For $\phi_{1}$ and $\phi_{2}$ the result follows from Proposition 2.9 by the reasoning in Remark 2.10. ∎
We now move on to the logistic loss function.
Example D.4.
Let $f(x)=\log(1+\exp(-\alpha^{\top}x))$. Then, $f$ is $(L,\bar{L})$-anisotropically smooth relative to $\phi_{i}$ for any $\bar{L}>0$ and $L>L_{\rm logistic}$ defined in the second row of Table 2.
Proof.
The gradient and the hessian of $f$ are given respectively by
@@
$a_{i}(\bar{L}^{-1}\nabla f(x))\frac{\|\alpha\|^{2}}{\exp(-\alpha^{\top}x)(1+\exp(\alpha^{\top}x))^{2}}<L\bar{L}.$ (39)
The results from Table 2 for $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$ then follow respectively from
$\frac{\|\alpha\|^{2}}{(1+\exp(-\alpha^{\top}x))\sqrt{\bar{L}^{2}(1+\exp(\alpha^{\top}x))^{2}}+\|\alpha\|^{2}}$ $\leq\frac{\|\alpha\|^{2}}{\sqrt{16\bar{L}^{2}+\|\alpha\|^{2}}}$ $<L,$
$\frac{\|\alpha\|^{2}}{(\bar{L}(1+\exp(\alpha^{\top}x))+\|\alpha\|)(1+\exp(-\alpha^{\top}x))}$ $\leq\frac{\|\alpha\|^{2}}{4\bar{L}+\|\alpha\|}$ $<L,$
$\frac{\bar{L}\|\alpha\|^{2}}{(\bar{L}(1+\exp(\alpha^{\top}x))+\|\alpha\|)^{2}\exp(-\alpha^{\top}x)}$ $\leq\frac{\bar{L}\|\alpha\|^{2}+\|\alpha\|^{3}}{4(\bar{L}+\|\alpha\|)^{2}}$ $<L.$
Note that in this case as well, $H\nabla^{2}f(x)$ is a symmetric matrix and it follows from Proposition D.2 that the operator $T_{{\delta L^{-1}},{\bar{L}^{-1}}}$ is injective for any $\delta<1$. The growth condition in Proposition 2.9 is satisfied automatically, since $f$ is bounded. ∎
D.2 Gradient clipping
@@
We now consider the case that $\phi=\frac{1}{2}\|\cdot\|^{2}+\delta_{\overline{\mathbb{B}}(0,1)}(\cdot)$. By (Themelis et al., 2019, p. 404), the generalized Jacobian of the preconditioner is given by
$\partial_{C}(\nabla\phi^{*})(y)=\begin{cases}\{I\},&\textnormal{if $\|y\|<1$,}\\ \operatorname{con}\{I,\Pi_{y^{\perp}}\},&\textnormal{if $\|y\|=1$,}\\ \{\|y\|^{-1}\Pi_{y^{\perp}}\},&\textnormal{if $\|y\|>1$.}\end{cases}$
where $\Pi_{y^{\perp}}$ denotes the projection matrix onto the orthogonal complement of the subspace spanned by $y$. The second-order condition (5) then becomes
@@
$\begin{cases}\lambda_{\rm max}(\nabla^{2}f(x))<L\bar{L},&\textnormal{if $\|\nabla f(x)\|<\bar{L}$,}\\ \lambda_{\rm max}\left(\alpha\nabla^{2}f(x)+(1-\alpha)\bar{L}\Pi_{\nabla f(x)^{\perp}}\nabla^{2}f(x)\right)<L\bar{L},&\textnormal{if $\|\nabla f(x)\|=\bar{L}$,}\\ \lambda_{\rm max}(\Pi_{\nabla f(x)^{\perp}}\nabla^{2}f(x))<L\|\nabla f(x)\|,&\textnormal{if $\|\nabla f(x)\|>\bar{L}$,}\end{cases}$
for all $\alpha\in[0,1]$. In particular, for $\alpha=1$, the second condition becomes $\nabla^{2}f(x)\prec L\bar{L}I$, and using the fact that $\lambda_{\rm max}(\Pi_{\nabla f(x)^{\perp}})=1$, we also have $\lambda_{\rm max}((\alpha I+(1-\alpha)\Pi_{\nabla f(x)^{\perp}})\nabla^{2}f(x))\leq\lambda_{\rm max}(\nabla^{2}f(x))$ such that this clause can be merged with the first case. To rewrite the last case in terms of symmetric matrices, we note that $\lambda_{\rm max}(\Pi_{\nabla f(x)^{\perp}}\nabla^{2}f(x))=\lambda_{\rm max}(\Pi_{\nabla f(x)^{\perp}}^{2}\nabla^{2}f(x))=\lambda_{\rm max}(\Pi_{\nabla f(x)^{\perp}}\nabla^{2}f(x)\Pi_{\nabla f(x)^{\perp}})$ where the last equality follows by (Horn & Johnson, 2012, Theorem 1.3.22). The second-order condition is therefore
$\begin{cases}\nabla^{2}f(x)\prec L\bar{L}I,&\textnormal{if $\|\nabla f(x)\|\leq\bar{L}$,}\\ \Pi_{\nabla f(x)^{\top}}\nabla^{2}f(x)\Pi_{\nabla f(x)^{\top}}\prec L\|\nabla f(x)\|I,&\textnormal{if $\|\nabla f(x)\|>\bar{L}$.}\end{cases}$
Note that the first case is akin to standard $L$-smoothness while the second case is reminiscent of $(L_{0},L_{1})$-smoothness with $L_{0}=0$ but restricted to some subspace. In particular, if $\nabla^{2}f(x)=C(x)\nabla f(x)\nabla f(x)^{\top}$ for some $C:\mathbb{R}^{n}\to\mathbb{R}$, then the second case is always satisfied, and we only require that $\nabla^{2}f(x)\prec L\bar{L}I$ for $\|\nabla f(x)\|\leq\bar{L}$. This is the case for both of the examples considered in Subsection D.1.